# Masked Mixers for Language Generation and Retrieval

### v2

@@ Abstract
Attention mechanisms that confer selective focus on a strict subset of input elements are nearly ubiquitous in language models today. We posit there to be downside to the use of attention: most input information is lost. In support of this idea we observe poor input representation accuracy in transformers and more accurate representation in what we term masked mixers, which replace self-attention with masked convolutions. The masked mixer learns causal language modeling more efficiently than early transformer implementations and even outperforms optimized, current transformers when training on small (<512) but not larger context windows. Evidence is presented for the hypothesis that differences in transformer and masked mixer training efficiencies for various tasks are best predicted by input representation accuracy, or equivalently global invertibility. We hypothesize that the information loss exhibited by transformers would be more detrimental to retrieval than generation, as the former is more closely approximated by a bijective and thus invertible function. We find that masked mixers are more effective retrieval models both when the pretrained embedding model is unchanged as well as when the embedding model is modified via cosine similarity-based InfoNCE loss minimization. A small masked mixer is shown to outperform a large and near state-of-the-art transformer-based retrieval model, despite the latter being trained with many orders of magnitude more data and compute.
@@ 1 Introduction
In spite of this impressive success, it is less clear how efficiently transformers or other models using attention may be trained for various language tasks: many types of architectures experience increases in goodness with an increased parameter numbers and dataset sizes, making early claims that the transformer scaled in superior ways to other models somewhat dubious. One way to investigate this question indirectly is to observe how information passes from the input to various hidden layers of the model, which can be considered to be closely related to the question of how these hidden layers represent the model’s input. Current investigations on this topic often focus on the method by which transformers transfer information to and from different tokens (c.f. [7]), but we focus on a simpler question: how much rather than how information is passed from the input to various parts of the model. In particular, we focus on the information present in the model’s last hidden layer, as this is most commonly used for tasks such as causal modeling (with a language modeling head applied) and retrieval.
@@
In this work we observe that transformer model hidden layers do not contain enough information to allow for accurate input reconstruction via gradient-based input representation algorithms, whereas the introduced ‘masked mixer’ models with self-attention substituted for masked convolutions do. We then compare causal language training efficiencies for a number of modeling tasks and find that the more a task requires something similar to what would be termed model invertibility, the more the masked mixer tends to outperform the transformer for both small and larger datasets. In general we corroborate the findings of the original vision-based mixers which that find that attention is superfluous to much of the current model performance observed [10; 11].
We investigate the performance of masked mixers and transformers for language retrieval task where a query to text matching task. We introduce a novel approach to training an accessory retrieval model using the embeddings of a CLM-pretrained model, and find that the masked mixer’s embeddings result in far more accurate retrieval. We then investigate the ability of masked mixers and transformers to perform retrieval when the embedding model is trained directly via InfoNCE-based cosine similarity loss, and find that once again masked mixers are substantially more accurate than transformers in this paradigm. We conclude with the observation that a masked mixer trained using these methods exhibits more accurate retrieval than a near state-of-the-art transformer model trained using the same techniques but with approximately 50,000x the pretraining compute to provide more context into the scale of the difference between masked mixer and transformer retrieval abilities.
@@ 1.1 Related Work
Elaborations on the use of convolutions in place of attention have met with success in attempts to model language with sub-quadratic complexity in the sequence dimension. In [14], Fu and colleagues replace attention with Monarch matrix-based sequence transformations for efficient sub-quadratic language modeling, but motivated by that goal the authors use a Monarch matrix-specific causal parametrization rather than parameter masking for causal language modeling. Poli and colleagues [15] introduced the Hyena architecture consisting of a recurrent structure with long convolutions combined with a gating operation with the same goal, and also do not perform direct masking. We comment that these works provide valuable insight into achieving near-linear complexity language modeling via efficient convolutional operations, but our approach to efficient language modeling focuses on global invertibility predicting training efficiency rather than sequence complexity and thus we proceed with an entirely different theoretical framework and do not investigate sub-quadratic complexity convolutional operations (Monarch or Toeplitz weight matrices etc.) which are by nature relatively inefficient to train with respect to loss achieved per compute quantity applied.
@@
To appreciate the motivation behind this work it is vital to understand that accurate input representation is equivalent to global invertibility only within the feasible input domain, not any arbitrary input. This is important to appreciate because it explains how models that are composed of non-invertible transformations (such as every linear layer in this work) can exhibit global invertibility, and is significant departure from work focused on invertible model transformations such a attention [17].
@@ 2 Our contribution
    The masked mixer, a modification of the MLP mixer for language modeling along with an inference method
@@
    An non-trivial autoencoding method for sequence models with excellent retrieval characteristics
@@
    A retrieval training method in which an accessory model is trained to match embeddings
@@
The primary insight behind this work is that differences in a model’s input representation accuracy, or equivalently its global invertibility, are capable of predicting a wide range of differences in training efficiency and performance between transformers and models of identical architecture except for the substitution of attention for linear transformations. Using this principle as a guide, we introduce a new foundational model architecture that far surpasses the performance of a modern and optimized transformer for retrieval.
@@ 3 Accurate self and non-self token representation in Masked Mixers
How much input information does a model contain in its hidden layers? More specifically, given some hidden layer values $O_{l}(a,\theta)$ can we regenerate the model’s input $a$ or in other words can we effectively invert the model? We use a gradient descent-based approach to address this question for language models. With respect to information that can be used to uniquely identify a language input, the answer for transformers is not very much. We introduce the masked mixer and show that this model accurately represents its inputs before and after training, and even accurately represent a limited number of non-self tokens.
@@ 3.1 Input Representation Background
In this work we measure the information present in a model’s hidden layer representation of the input by attempting to recover the input using gradient descent on an initially random input, minimizing a chosen metric on the activations of that hidden layer as described in [18]. The goal is to invert a model such that information in the hidden layer’s activations representation is used to identify the input that resulted in those activations.
@@
We use a scheduled learning rate $\eta$ that decreases linearly from $\eta$ to $\eta/10$ as $n\to N$ (ie the iteration number $n$ is increased until it reaches the final number of iterations $N$, which we set somewhat arbitrarily to $N=500$) which empirically results in the fastest optimization with the fewest number of steps taken $N$.
@@
To be specific, we optimize the embedding of the input rather than the input itself. The embedding may be computed by matrix-vector multiplication using the embedding weight matrix $W$ such that $e=Wa$. With an initially random embedding $e_{0}=\mathcal{N}(e,\mu=1/2,\sigma=1/20)$ we perform gradient descent on the embedding using an $L^{1}$ metric on the layer in question as denoted in (2).
@@
In general the embedding weight matrix $W$ is non-square and non-invertible. $e_{N}$ is therefore converted back to the generated input $a_{N}$ via left multiplication by the Moore-Penrose pseudo-inverse $W^{+}$ via (3), which is a form of generalized matrix inverse defined by Equation (4).
@@
If positional encoding is applied to the embedding, it is subtracted before (3) is applied. This process is illustrated in Figure 1 for convenience.
@@
If the gradient is effectively backpropegated to the input $e_{n}$ in Equation (2) then $||O_{l}(e_{N},\theta)-O_{l}(e,\theta)||_{1}<\epsilon$ for some small $\epsilon$. We check this condition by adding a very small amount of Gaussian noise to the initial input embedding before observing the metric distance of the output of this noised input to $a$, ie $||O_{l}(e+{N}(e,\mu=0,\sigma=1/20),\theta))-O_{l}(e,\theta)||_{1}=\epsilon$. This allows us to confirm that iterations of (2) results in sufficient minimization of the metric on the output (in this case $L^{1}$) such that the corresponding $a$ and $a_{N}$ are nearly equivalent for the model. The particular amount of Gaussian noise to add in order to specify $\epsilon$ was determined in part by observing that the chosen value leads to no token changes upon pseudoinversion, meaning that typically $W^{+}e=W^{+}(e+\epsilon)$ although this depends somewhat on the transformation via the word-token embedding weights $W$.
@@ 3.2 Masked mixer architecture
Elsewhere it was observed that large and otherwise capable language models (Llama-2 7b and 70b etc.) exhibit relatively inaccurate input representation for all but the shallowest few hidden layers [19]. It was also observed that vision MLP mixer models ([11; 10]) have superior input representation accuracy for non-self tokens compared to vision transformers [20], leading us to wonder if a mixer adapted to causal language modeling could retain this accurate input representation. We hypothesized that adapting an MLP-mixer to the process of language generation would give us a model suitable for autoregressive language generation and other tasks, but with much more accurate input representation abilities.
@@
where ‘f’ and ‘d’ are both equal to the number of tokens in the context window and ‘k’ is the 1-dimensional kernel size. A 1D convolution with a kernel size of 1 can also be implemented using ‘torch.linear’ on the sequence dimension in which the triangular weight masking is a more pithy ‘conv.weight.data = torch.tril(conv.weight)’ after reshaping, but experimentally this does not exhibit the throughput as well as the 1D convolutional kernel method above nor is it as flexible.
@@ 3.3 Masked mixers but not transformers exhibit accurate input representations
We define this metric as follows: given input $x$ and generated representation of that input $y$ such that each element of $x$ is an integer corresponding to an element in the tokenizer set $\{t_{m}\}$, which may be stated precisely as
@@
then the normalized Hamming metric is given in Equation (6). In words, this metric is the fraction of indices of the input where the generated representation’s token does not match the input’s token while ignoring the input’s pad token positions. The smaller the normalized Hamming distance, the more similar the input is to the model’s representation and the more information that representation carries. We measure before training as well as during and after 12 hours of RTX 3060 training on the first 2M examples of the TinyStories [23] dataset.
@@
As observed elsewhere for larger language models, small transformers designed for TinyStories modeling exhibit very poor input representation such that the information present in the last hidden layer is incapable of recovering very little of an input when (2) is applied (Figure 3), although this improves somewhat with training. Masked mixers exhibit near-perfect representation before training, and larger models retain this characteristic even after training on Tinystories (Figure 3).
Figure 3: Masked mixers exhibit more accurate input representation than transformers. All models are $n_{l}=8$ and all transformers are llama models with the default $n_{h}=32$. In d) the transformer is $d_{m}=256$ and mixer $d_{m}=512$.
Masked mixers also exhibit much more accurate input representation than transformers when gradient descent is performed using only the last hidden layer of the last token (Figure 3b), indicating that they pass more information between tokens than transformers. Neither transformers nor masked mixers exhibit accurate non-self token representation for larger context windows ($c=512$ etc.) if only the last token’s last hidden layer is used for backpropegation, indicating that there are limits to the inter-token information bandwidth in mixers as well as transformers.
The original MLP-mixer architecture employs two sequential convolutional operations between each input token, with an expansion factor (usually set to two) for the layer between convolutions. We find that models with only one convolution operation between tokens (a ‘flat’ masked mixer) have superior inter-token information transfer (Figure S3), but that their self-token representation power is very similar.
Masked mixers of a sufficient $d_{m}$ are in some sense biased towards accurate input representation, as untrained models nearly always exhibit perfect input representation regardless of their actual randomly initialized values. In this vein, transformers of a sufficient $d_{model}$ are biased against accurate input representation. It is interesting therefore to observe that masked mixer and transformer input representation abilities converge somewhat during training (Figure 3d), which suggests that for the TinyStories dataset there is some range of optimal input representation accuracies that is associated with minimization of all-next-token prediction during causal language modeling. This convergence is very slow, however, especially for larger models such that the model type’s initial bias determines most behavior after training as well (Figure 3c).
@@ 4 Masked Mixer and Transformer Causal Language Modeling Efficiency
The ability of large transformer models to provide useful language modeling without accurate input representation suggests that accurate input representation is not necessary for many of the abilities of that model type. We reasoned that although input representational accuracy does not necessarily provide greater language modeling capabilities, it could perhaps lead to more efficient training.
To test the training efficiency of the Masked Mixer compared to the Transformer, we employ the same two datasets as for representation studies: the first train using 2M training examples of TinyStories with all evaluation examples with a Llama-2 tokenizer of size 4096 trained on that dataset, where models are trained using a correspondingly small fixed compute budget of 12 hours on an Nvidia RTX 3060, or 2.25 hours on a 4x Nvidia V100 cluster. We then train on a much larger and more varied dataset, specifically a 10 billion token subset of a extensively refined of recent Common Crawl snapshots called the FineWeb-edu [24] (hereafter termed ‘FineWeb’ for simplicity) with approximately ten times the compute amount, or approximately 20 hours on a 4x V100 server, with 200k steps taken per training run.
As the Llama -style transformer uses the same Cross-Entropy Loss function (with loss masking on pad token outputs) on the language modeling head as the Masked Mixer with the same hyperparameters (a context window of 512 unless otherwise noted, AdamW optimizer [25], etc.), we can directly compare the loss achieved after the fixed compute quantity has been applied.

### v3

@@ Abstract
Attention mechanisms that confer selective focus on a strict subset of input elements are nearly ubiquitous in language models today. We posit there to be downside to the use of attention: most input information is lost. In support of this idea we observe poor input representation accuracy in transformers and more accurate representation in what we term masked mixers, which replace self-attention with masked convolutions. The masked mixer learns causal language modeling more efficiently than early transformer implementations and even outperforms optimized, current transformers when training on small ($n_{ctx}<512$) but not larger context windows. Evidence is presented for the hypothesis that differences in transformer and masked mixer training efficiencies for various tasks are best predicted by input representation accuracy, or equivalently global invertibility. We hypothesize that the information loss exhibited by transformers would be more detrimental to retrieval than generation, as the former is more closely approximated by a bijective and thus invertible function. We find that masked mixers are more effective retrieval models both when the pretrained embedding model is unchanged as well as when the embedding model is modified via cosine similarity-based InfoNCE loss minimization. A small masked mixer is shown to outperform a large and near state-of-the-art transformer-based retrieval model, despite the latter being trained with many orders of magnitude more data and compute.
@@ 1 Introduction
In spite of this impressive success, it is less clear how efficiently transformers or other models using attention may be trained for various language tasks: many types of architectures experience increases in goodness with an increased parameter numbers and dataset sizes, making early claims that the transformer scaled in superior ways to other models somewhat dubious. One way to investigate this question indirectly is to observe how information passes from the input to various hidden layers of the model, which can be considered to be closely related to the question of how these hidden layers represent the model’s input. Current investigations on this topic often focus on the method by which transformers transfer information to and from different tokens (cf. [7]), but we focus on a simpler question: how much rather than how information is passed from the input to various parts of the model. In particular, we focus on the information present in the model’s last hidden layer, as this is most commonly used for tasks such as causal modeling (with a language modeling head applied) and retrieval.
@@
In this work we observe that transformer model hidden layers do not contain sufficient uniquely identifying information to allow for accurate input reconstruction via gradient-based input representation algorithms, whereas the introduced ‘masked mixer’ models with self-attention substituted for masked convolutions do and are in that sense globally invertible. We then compare causal language training efficiencies for a number of modeling tasks and find that the more a task requires something similar to what would be termed model invertibility, the more the masked mixer tends to outperform the transformer for both small and larger datasets. In general we corroborate the observations of original vision-based MLP mixer work in finding that attention is superfluous to much of the transformer’s performance [10; 11].
We investigate the performance of masked mixers and transformers for language retrieval task where a query to text matching task. We introduce a novel approach to training an accessory retrieval model using the embeddings of a CLM-pretrained model, and find that the masked mixer’s embeddings result in far more accurate retrieval. We then investigate the ability of masked mixers and transformers to perform retrieval when the embedding model is trained directly via InfoNCE-based cosine similarity loss, and find that once again masked mixers are substantially more accurate than transformers in this paradigm. We conclude with the observations that a masked mixer exhibits more accurate retrieval than a large and near state-of-the-art transformer-based retrieval model pretrained with approximately 10,000x the compute, as well as scaling data to show that transformers require many orders of magnitude more retrieval data to match the masked mixer’s accuracy.
@@ 1.1 Related Work
Elaborations on the use of convolutions in place of attention have met with success in attempts to model language with sub-quadratic complexity in the sequence dimension. In [14], Fu and colleagues replace attention with Monarch matrix-based sequence transformations for efficient sub-quadratic language modeling, but motivated by that goal the authors use a Monarch matrix-specific causal parametrization rather than parameter masking for causal language modeling. Poli and colleagues [15] introduced the Hyena architecture consisting of a recurrent structure with long convolutions combined with a gating operation with the same goal, and also do not perform direct masking. We comment that these works provide valuable insight into achieving near-linear complexity language modeling via efficient convolutional operations, but our approach to efficient language modeling focuses on global invertibility predicting training efficiency rather than sequence complexity and thus we proceed with an entirely different theoretical framework and do not investigate sub-quadratic complexity convolutional operations (Monarch or Toeplitz weight matrices etc.) which trade inference efficiency for training inefficiencies in terms of loss achieved per compute applied.
@@
To appreciate the motivation behind this work it is vital to understand that accurate input representation is equivalent to global invertibility only within the feasible input domain, not any arbitrary input. This is important to appreciate because it explains how models that are composed of non-invertible transformations (such as every linear layer in this work) can exhibit global invertibility, and is significant departure from work focused on converting model transformations to invertible versions of those transformations [17].
@@ 2 Our Contribution
    The masked mixer, a modification of the MLP mixer for language modeling
@@
    An non-trivial autoencoding method for sequence models with promising retrieval characteristics
@@
    A high-throughput retrieval sampling method and architecture in which a model is trained to match embeddings
@@
The primary insight behind this work is that differences in a model’s input representation accuracy, or equivalently its global invertibility, are capable of predicting a wide range of differences in training efficiency and performance between transformers and models of identical architecture except for the substitution of attention for linear transformations. Using this principle as a guide, we introduce a new foundational model architecture that far surpasses the performance of the transformer for retrieval and autoencoding.
@@ 3 Accurate Self and Non-Self Token Representation in Masked Mixers
How much input information does a model contain in its hidden layers? More specifically, given some hidden layer values $O_{l}(a,\theta)$ can we regenerate the model’s input $a$ or in other words can we globally invert the model? We use a gradient descent-based approach to address this question for language models. With respect to information that can be used to uniquely identify a language input, the answer for transformers is not very much. We introduce the masked mixer and show that this model accurately represents its inputs before and after training, and even accurately represent a limited number of non-self tokens.
@@ 3.1 Input representation background
In this work we measure the information present in a model’s hidden layer representation of the input by attempting to recover the input using gradient descent on an initially random input, minimizing a chosen metric on the activations of that hidden layer as described in [18]. The goal is to invert a model such that information in the hidden layer’s activations (which may be thought of as that layer’s representation) is used to identify the model’s input.
@@
We use a scheduled learning rate $\eta$ that decreases linearly from $\eta$ to $\eta/10$ as $n\to N$ (where the iteration number $n$ is increased until it reaches the final number of iterations $N$, which we set somewhat arbitrarily to $N=500$) which empirically results in the fast optimization with a reasonable $N$.
@@
To be specific, we optimize the embeddings of the input tokens rather than the input itself. The embeddings may be computed by matrix-vector multiplication using the embedding weight matrix $W$ such that $e=Wa$. With an initially random embedding $e_{0}=\mathcal{N}(e,\mu=1/2,\sigma=1/20)$ we perform gradient descent on the embedding using an $L^{1}$ metric on the layer in question as denoted in (2).
@@
In general the embedding weight matrix $W$ is non-square and non-invertible. $e_{N}$ is therefore converted back to the generated input $a_{N}$ via left multiplication by the Moore-Penrose pseudo-inverse $W^{+}$ via (3), which is a form of generalized matrix inverse defined by Equation (4). If positional encoding is applied to the embedding, it is subtracted before (3) is applied. This process is illustrated in Figure 1 for convenience.
@@
If the gradient is effectively backpropegated to the input $e_{n}$ in Equation (2) then $||O_{l}(e_{N},\theta)-O_{l}(e,\theta)||_{1}<\epsilon$ for some small $\epsilon$. We establish the value of $\epsilon$ in order to check this condition by adding a very small amount of Gaussian noise to the initial input embedding before observing the metric distance of the output of this noised input to $a$ such that $\epsilon=||O_{l}(e+{N}(e,\mu=0,\sigma=1/20),\theta))-O_{l}(e,\theta)||_{1}$. This check allows us to confirm that iterations of (2) results in sufficient minimization of the metric on the output (in this case $L^{1}$) such that the corresponding $a$ and $a_{N}$ are nearly equivalent for the model. The particular amount of Gaussian noise to add in order to specify the output shift $\epsilon$ was determined in part by observing that the chosen value leads to no token changes upon pseudoinversion, meaning that typically $W^{+}e=W^{+}(e+{N}(e,\mu=0,\sigma=1/20))$ although this depends somewhat on the transformation via the word-token embedding weights $W$.
@@ 3.2 Masked mixer architecture
Elsewhere it was observed that large and otherwise capable language models (Llama-2 7b and 70b etc.) exhibit relatively inaccurate input representation for all but the shallowest few hidden layers [19]. It was also observed that vision MLP mixer models ([11; 10]) have superior input representation accuracy for non-self tokens compared to vision transformers [20], leading us to wonder if a mixer adapted to causal language modeling could exhibit this accurate input representation. We hypothesized that adapting an MLP-mixer to the process of language generation would give us a model suitable for autoregressive language generation and other tasks, but with much more accurate input representation abilities.
@@
where ‘f’ and ‘d’ are both equal to the number of tokens in the context window and ‘k’ is the 1-dimensional kernel size. A 1D convolution with a kernel size of 1 can also be implemented using ‘torch.linear’ on the sequence dimension in which the triangular weight masking is a more pithy ‘conv.weight.data = torch.tril(conv.weight)’ after reshaping, but experimentally this does not exhibit as much throughput as the 1D convolutional kernel method above nor is it as flexible for non-unitary kernels.
@@ 3.3 Masked mixers but not transformers exhibit accurate input representations
We define this metric as follows: given input $x$ and generated representation of that input $y$ such that each element of $x$ is an integer corresponding to an element in the tokenizer set $\{t_{m}\}$, which may be stated precisely as shown in Equation 5, then the normalized Hamming metric is given in Equation (6). In words, this metric is the fraction of indices of the input where the generated representation’s token does not match the input’s token while ignoring the input’s pad token positions. The smaller the normalized Hamming distance, the more similar the input is to the model’s representation and the more information that representation carries. We measure before training as well as during and after 12 hours of RTX 3060 training on the first 2M examples of the TinyStories [23] dataset.
@@
As observed elsewhere for larger language models, small transformers designed for TinyStories modeling exhibit very poor input representation such that the information present in the last hidden layer is capable of recovering little if any of an input when (2) is applied (Figure 3), although this improves somewhat with training. Masked mixers exhibit near-perfect representation before training, and larger models retain this characteristic even after training on Tinystories (Figure 3).
Figure 3: Masked mixers exhibit more accurate input representation than transformers. All models are $n_{l}=8$ and all transformers are llama architectures with $n_{h}=32$. In d) the transformer is $d_{m}=256$ and mixer $d_{m}=512$, trained on TinyStories.
Masked mixers also exhibit much more accurate input representation than transformers when gradient descent is performed using only the last hidden layer of the last token (Figure 3b), indicating that they pass more information between tokens than transformers. Neither transformers nor masked mixers exhibit accurate non-self token representation for larger context windows ($c=512$ etc.) if only the last token’s last hidden layer is used for backpropagation, indicating that there are limits to the inter-token information bandwidth in mixers as well as transformers before training.
The original MLP-mixer architecture employs two sequential convolutional operations between each input token, with an expansion factor (usually set to two) for the layer between convolutions. We find that models with only one convolution operation between tokens (a ‘flat’ masked mixer) have superior inter-token information transfer (Figure S1), but that their self-token representation power is similar.
Masked mixers of a sufficient $d_{model}$ are in some sense biased towards accurate input representation as untrained models nearly always exhibit perfect input representation regardless of their actual randomly initialized values. In this vein, transformers of a sufficient $d_{model}$ are biased against accurate input representation. It is interesting therefore to observe that masked mixer and transformer input representation abilities converge somewhat during training (Figure 3d), which suggests that for the TinyStories dataset there is some range of optimal input representation accuracies that is associated with minimization of all-next-token prediction during causal language modeling. This convergence is very slow, however, especially for larger models such that the model type’s initial bias determines most behavior after training as well (Figure 3c).
@@ 4 Masked Mixer and Transformer Causal Language Modeling Efficiency
The ability of large transformer models to provide useful causal language modeling without accurate input representation suggests that accurate input representation is not necessary for that task. We reasoned that although input representational accuracy does not necessarily provide greater language modeling capabilities, it could perhaps lead to more efficient training in certain cases.
To begin to test the training efficiency of the Masked Mixer compared to the Transformer, we employ the same two datasets as for representation studies: the first train using 2M training examples of TinyStories with all evaluation examples with a Llama-2 tokenizer of size 4096 trained on that dataset, where models are trained using a correspondingly small fixed compute budget of 12 hours on an Nvidia RTX 3060, or 2.25 hours on a 4x Nvidia V100 server. We then train on a much larger and more varied dataset, specifically a 10 billion token subset of a extensively refined of recent Common Crawl snapshots called the FineWeb-edu [24] (hereafter termed ‘FineWeb’ for simplicity) with ten times the compute amount, or approximately 20 hours on a 4x V100 server, with 200k steps taken per training run.
As the Llama -style transformer uses the same Cross-Entropy Loss function (with loss masking on pad token outputs) on the language modeling head as the Masked Mixer with the same hyperparameters (a context window of 512 unless otherwise noted, AdamW optimizer [25], etc.), we can directly compare the loss achieved after the fixed compute quantity has been applied, around $94\times 10^{1}$ 9 Tensor FLOPs for a 20 hour training run on the 4x V100 server.